OpenStudy (anonymous):

major help on this question a. Calculate the scalar projection of (5,3) onto (7,-2). b. Interpret this projection graphically.

OpenStudy (anonymous):

this is what i have so far\[(5\times7-3\times2)/\sqrt{53}=29/\sqrt{53}\]

OpenStudy (anonymous):

\[a \times b/|a|<A>=29/53<7,-2>\]